A wave represented by y= (3.00m) sin (2.5t-3.14x). Find the (a) amplitude, (b) angular frequency, (c) frequency, (d) wave number, (e) wavelength, and (f) speed of the wave


1
Expert's answer
2021-04-20T16:49:45-0400

Answer

Amplitude

A=3.00m

Angular frequency =2.5rad/s

Frequency =2.5/2(3.14) =0.398Hz

Wave number k=3.14 "m^{-1}"


Wavelength

"\lambda=\frac{2 \pi}{k}\\=\lambda=\frac{2 \pi}{3.14}=2m"


"v=f\lambda=0.398(2) =0.79m/s"